# Statistical Inference for Pairwise Comparison Models

### v2

@@ B.2 Proof of Lemma 15
Under Assumption 5, $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})\}_{j:(i,j)\in{\mathcal{E}}}$ are subgaussian random variables with subgaussian norms uniformly bounded by $c_{n,2}$. According to (Vershynin, 2018, Lemma 2.7.6), $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}\}_{j:(i,j)\in{\mathcal{E}}}$ are sub-exponential random variables with sub-exponential norms uniformly bounded by $c_{n,2}^{2}$. It follows from (Vershynin, 2018, Proposition 2.7.1, Exercise 2.7.10) that
@@
$\|\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\|_{2}$$\lesssim\left\{(c_{n,4}\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}^{2}+c_{n,2}^{2})nq_{n}\right\}^{1/2}\lesssim c_{n,2}(nq_{n})^{1/2},$ (31)
@@
$\lambda_{2}\{\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}\gtrsim c_{n,3}np_{n}\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2},$ (32)
where the extra term $c_{n,3}np_{n}$ stems from a lower bound for the singular value of the normalization matrix. Combining (30)–(32), we conclude
$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\lesssim\frac{c_{n,2}c_{n,4}^{2}q_{n}^{5/2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (33)
Putting (27), (28), (29), and (33) together,
[…]
$\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}^{3}q_{n}^{5/2}(\log n)^{2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (34)
Consequently, combining (25) and (34), we have
@@
Appendix C Proofs of Corollary 10 and Theorems 11-12
@@ C.1 Proof of Corollary 10
Proof  If $M^{*}<\infty$, then $\sup_{n}c_{n,1}<1$ so that Assumption 4 holds under condition $(\log n)^{8}/(np_{n})\to 0$. Since ${\mathbb{A}}$ is finite and $g(x;y)$ is continuous in $y$, $g(x;y)$ is uniformly bounded over ${\mathbb{A}}\times[-M^{*},M^{*}]$, implying $\sup_{n}c_{n,2}<\infty$. Additionally, by the strict log-concavity of $f$, $\partial_{2}g(x;y)>0$, which combined with the continuity of $\partial_{2}g(x;y)$, the boundedness of $M^{*}$, and the finiteness of ${\mathbb{A}}$ implies $0<\inf_{n}c_{n,3}\leq\sup_{n}c_{n,4}<\infty$. A similar argument shows $\sup_{n}c_{n,5}<\infty$. The proof is finished by appealing to (3).
@@ C.2 Proof of Theorem 11
$\{{\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})\}_{k}$ =$=-\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{k}+o_{p}(1),$
@@ C.3 Proof of Theorem 12
According to Slutsky’s Theorem and Thereom 8, it suffices to show that for any fixed $i\in[n]$ that
@@
where the last step follows assuming (3) holds. Meanwhile,

### v3

@@ B.2 Proof of Lemma 3
Under Assumption 2, $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})\}_{j:(i,j)\in{\mathcal{E}}}$ are subgaussian random variables with subgaussian norms uniformly bounded by $c_{n,2}$. According to (vershynin2018high, Lemma 2.7.6), $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}\}_{j:(i,j)\in{\mathcal{E}}}$ are sub-exponential random variables with sub-exponential norms uniformly bounded by $c_{n,2}^{2}$. It follows from (vershynin2018high, Proposition 2.7.1, Exercise 2.7.10) that
@@
$\|\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\|_{2}$ $\lesssim\left\{(c_{n,4}\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}^{2}+c_{n,2}^{2})nq_{n}\right\}^{1/2}\lesssim c_{n,2}(nq_{n})^{1/2},$ (B.20)
@@
$\lambda_{2}\{\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}\gtrsim c_{n,3}np_{n}\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2},$ (B.21)
where the extra term $c_{n,3}np_{n}$ stems from a lower bound for the singular value of the normalization matrix. Combining (B.19)–(B.21), we conclude
$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\lesssim\frac{c_{n,2}c_{n,4}^{2}q_{n}^{5/2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (B.22)
Putting (B.16), (B.17), (B.18), and (B.22) together,
$(A.1)$ $\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\left(\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}+\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\right)$
$\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}^{3}q_{n}^{5/2}(\log n)^{2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (B.23)
Consequently, combining (B.14) and (B.23), we have
@@
Appendix C Proofs of Corollary 1 and Theorems 2-3
@@ C.1 Proof of Corollary 1
Proof.
If $M^{*}<\infty$, then $\sup_{n}c_{n,1}<1$ so that Assumption 1 holds under condition $(\log n)^{8}/(np_{n})\to 0$. Since $\mathbb{A}$} is finite and $g(x;y)$ is continuous in $y$, $g(x;y)$ is uniformly bounded over ${\mathbb{A}}\times[-M^{*},M^{*}]$, implying $\sup_{n}c_{n,2}<\infty$. Additionally, by the strict log-concavity of $f$, $\partial_{2}g(x;y)>0$, which combined with the continuity of $\partial_{2}g(x;y)$, the boundedness of $M^{*}$, and the finiteness of $\mathbb{A}$} implies $0<\inf_{n}c_{n,3}\leq\sup_{n}c_{n,4}<\infty$. A similar argument shows $\sup_{n}c_{n,5}<\infty$. The proof is finished by appealing to (3.3). $\Box$
@@ C.2 Proof of Theorem 2
$\{{\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})\}_{k}$ $=-\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{k}+o_{p}(1),$
@@ C.3 Proof of Theorem 3
According to Slutsky’s theorem and Theorem 1, it suffices to show that for any fixed $i\in[n]$ that
@@
where the last step follows assuming (3.3) holds. Meanwhile,
@@
Funding
R. Han was supported by the Hong Kong Research Grants Council (No. 14301821) and the Hong Kong Polytechnic University (P0044617, P0045351).
Relevant code
The code to reproduce the main results in this paper is available at https://github.com/RJ-HAN-STAT/Ranking_Code.